# Electric Dipole Moments in 5+3 Flavor Weak Effective Theory

### v1

@@ 4.4.4 ${C_{\!\underset{prrp}{ee}}^{{\tiny V,LR}}}$ onto dipole
$[C^{S1,RR}_{dd}]_{3333}$ $4\cdot 10^{-8}$ $8\cdot 10^{-8}$ $2\cdot 10^{-6}$ $-10\cdot 10^{-5}$ $-4\cdot 10^{-1}$ $1\cdot 10^{-2}$ $2\cdot 10^{-5}$ $-2\cdot 10^{-7}$
$[C^{S8,RR}_{dd}]_{3333}$ $-2\cdot 10^{-8}$ $-3\cdot 10^{-7}$ $2\cdot 10^{-5}$ $9\cdot 10^{-2}$ $2\cdot 10^{-2}$ $-4\cdot 10^{-6}$ $4\cdot 10^{-8}$
$[C^{S1,RR}_{dd}]_{1133}$ $5\cdot 10^{-2}$ $4\cdot 10^{-4}$ $5\cdot 10^{-5}$ $4\cdot 10^{-7}$
$[C^{S8,RR}_{dd}]_{1133}$ $2\cdot 10^{-2}$ $-2\cdot 10^{-3}$ $2\cdot 10^{-5}$ $-2\cdot 10^{-6}$
$[C^{S1,RR}_{dd}]_{2233}$ $5\cdot 10^{-2}$ $4\cdot 10^{-4}$ $2\cdot 10^{-7}$ $1\cdot 10^{-3}$ $8\cdot 10^{-6}$ $-3\cdot 10^{-8}$
$[C^{S8,RR}_{dd}]_{2233}$ $2\cdot 10^{-2}$ $-2\cdot 10^{-3}$ $6\cdot 10^{-8}$ $3\cdot 10^{-4}$ $-4\cdot 10^{-5}$ $-1\cdot 10^{-8}$
$[C^{S1,RR}_{dd}]_{1331}$ $-3\cdot 10^{-1}$ $7\cdot 10^{-3}$ $-6\cdot 10^{-8}$ $-3\cdot 10^{-4}$ $7\cdot 10^{-6}$ $1\cdot 10^{-8}$
$[C^{S8,RR}_{dd}]_{1331}$ $3\cdot 10^{-2}$ $1\cdot 10^{-2}$ $3\cdot 10^{-5}$ $1\cdot 10^{-5}$
$[C^{S1,RR}_{dd}]_{2332}$ $-3\cdot 10^{-1}$ $7\cdot 10^{-3}$ $-1\cdot 10^{-6}$ $-5\cdot 10^{-3}$ $1\cdot 10^{-4}$ $2\cdot 10^{-7}$
$[C^{S8,RR}_{dd}]_{2332}$ $3\cdot 10^{-2}$ $1\cdot 10^{-2}$ $1\cdot 10^{-7}$ $6\cdot 10^{-4}$ $2\cdot 10^{-4}$ $-3\cdot 10^{-8}$
$[C^{S1,RR}_{ud}]_{1133}$ $5\cdot 10^{-2}$ $-9\cdot 10^{-4}$ $2\cdot 10^{-5}$ $2\cdot 10^{-7}$
$[C^{S8,RR}_{dd}]_{1133}$ $2\cdot 10^{-2}$ $-2\cdot 10^{-3}$ $2\cdot 10^{-5}$ $-2\cdot 10^{-6}$
$[C^{S1,RR}_{uddu}]_{1331}$ $-3\cdot 10^{-1}$ $2\cdot 10^{-2}$ $-3\cdot 10^{-8}$ $-1\cdot 10^{-4}$ $-1\cdot 10^{-5}$
$[C^{S8,RR}_{uddu}]_{1331}$ $3\cdot 10^{-2}$ $9\cdot 10^{-3}$ $1\cdot 10^{-5}$ $-9\cdot 10^{-6}$
$[C^{S1,RR}_{ud}]_{2233}$ $-3\cdot 10^{-8}$ $3\cdot 10^{-4}$ $1\cdot 10^{-2}$ $8\cdot 10^{-5}$ $5\cdot 10^{-2}$ $-9\cdot 10^{-4}$
$[C^{S8,RR}_{ud}]_{2233}$ $-1\cdot 10^{-8}$ $8\cdot 10^{-5}$ $4\cdot 10^{-3}$ $1\cdot 10^{-3}$ $2\cdot 10^{-2}$ $-2\cdot 10^{-3}$
$[C^{S1,RR}_{uddu}]_{2332}$ $1\cdot 10^{-8}$ $2\cdot 10^{-7}$ $-1\cdot 10^{-3}$ $-6\cdot 10^{-2}$ $-6\cdot 10^{-3}$ $-3\cdot 10^{-1}$ $2\cdot 10^{-2}$
$[C^{S8,RR}_{uddu}]_{2332}$ $-3\cdot 10^{-8}$ $1\cdot 10^{-4}$ $7\cdot 10^{-3}$ $-5\cdot 10^{-3}$ $3\cdot 10^{-2}$ $9\cdot 10^{-3}$
$[C^{V1,LR}_{uddu}]_{1331}$ $-7\cdot 10^{-3}$ $-6\cdot 10^{-4}$ $-3\cdot 10^{-6}$ $-1\cdot 10^{-6}$
$[C^{V8,LR}_{uddu}]_{1331}$ $4\cdot 10^{-3}$ $5\cdot 10^{-3}$ $2\cdot 10^{-6}$ $-4\cdot 10^{-6}$
$[C^{V1,LR}_{dd}]_{1331}$ $-5\cdot 10^{-3}$ $3\cdot 10^{-3}$ $-5\cdot 10^{-6}$ $3\cdot 10^{-6}$
$[C^{V8,LR}_{dd}]_{1331}$ $4\cdot 10^{-3}$ $4\cdot 10^{-3}$ $4\cdot 10^{-6}$ $4\cdot 10^{-6}$
$[C^{V1,LR}_{dd}]_{2332}$ $-5\cdot 10^{-3}$ $3\cdot 10^{-3}$ $-3\cdot 10^{-8}$ $-9\cdot 10^{-5}$ $6\cdot 10^{-5}$
$[C^{V8,LR}_{dd}]_{2332}$ $4\cdot 10^{-3}$ $4\cdot 10^{-3}$ $3\cdot 10^{-8}$ $8\cdot 10^{-5}$ $8\cdot 10^{-5}$
$[C^{V1,LR}_{uddu}]_{2332}$ $-4\cdot 10^{-5}$ $-1\cdot 10^{-3}$ $-7\cdot 10^{-4}$ $-7\cdot 10^{-3}$ $-7\cdot 10^{-4}$
$[C^{V8,LR}_{uddu}]_{2332}$ $2\cdot 10^{-5}$ $10\cdot 10^{-4}$ $-2\cdot 10^{-3}$ $4\cdot 10^{-3}$ $5\cdot 10^{-3}$
$[C_{dG}]_{33}$ $-2\cdot 10^{-7}$ $-4\cdot 10^{-7}$ $-9\cdot 10^{-6}$ $-5\cdot 10^{-8}$ $5\cdot 10^{-4}$ $1\cdot 10^{0}$ $4\cdot 10^{-2}$ $-1\cdot 10^{-4}$ $1\cdot 10^{-6}$
$[C_{d\gamma}]_{33}$ $-2\cdot 10^{-7}$ $9\cdot 10^{-6}$ $3\cdot 10^{-2}$ $9\cdot 10^{-1}$ $-2\cdot 10^{-6}$ $2\cdot 10^{-8}$
Table 5: At the 2-loop level, the operator mixing of the bottom quark operators in $5+3$ WET (1st column) onto dipole operators (2nd row) in $3+1$ and $5+3$ flavor WET is shown. Here the entries represent the quantity $\eta_{i}^{j}(\mu_{\rm low},\mu_{\rm ew})$ Entries below $10^{-8}$ are dropped.
4.5 Numerical solution for the RGEs
As we have learned in the previous subsections, 1-loop and 2-loop RG running effects and threshold effects cause operator mixing leading to the expansion of operator basis for the EDMs. In this section, we provide the numerical solution to the RGEs, which allows to quantify the impact of such effects on the low-energy operators that contribute to EDMs at tree level. In Tabs. 4, 5 and 6, we show the numerical values of the quantity
for $\mu_{low}=2$GeV and $\mu_{ew}=91$GeV.
All semileptonic operators in 5+3 WET
$\eta_{i}^{j}(\mu_{\rm low},\mu_{\rm ew})$
$[C^{S,RR}_{eu}]_{1111}$ $6\cdot 10^{-8}$ $-3\cdot 10^{-7}$
$[C^{S,RR}_{ed}]_{1111}$ $-3\cdot 10^{-8}$ $-2\cdot 10^{-7}$
$[C^{S,RR}_{ed}]_{1122}$ $-3\cdot 10^{-8}$ $-3\cdot 10^{-6}$
@@ 4.5 Numerical solution for the RGEs
$[C^{T,RR}_{eu}]_{1111}$ $1\cdot 10^{-6}$ $-3\cdot 10^{-5}$ $1\cdot 10^{-4}$
$[C^{T,RR}_{ed}]_{1111}$ $-6\cdot 10^{-7}$ $-3\cdot 10^{-5}$ $-1\cdot 10^{-4}$
$[C^{T,RR}_{ed}]_{1122}$ $-6\cdot 10^{-7}$ $-3\cdot 10^{-5}$ $-3\cdot 10^{-3}$
$[C^{T,RR}_{eu}]_{1122}$ $7\cdot 10^{-2}$ $1\cdot 10^{-6}$ $-3\cdot 10^{-5}$
$[C^{S,RR}_{eu}]_{1122}$ $-2\cdot 10^{-4}$ $6\cdot 10^{-8}$
$[C^{T,RR}_{ed}]_{1133}$ $-1\cdot 10^{-1}$ $-6\cdot 10^{-7}$ $-3\cdot 10^{-5}$
$[C^{T,RR}_{ed}]_{2211}$ $-1\cdot 10^{-4}$ $-6\cdot 10^{-3}$
$[C^{T,RR}_{ed}]_{2222}$ $-1\cdot 10^{-4}$ $-6\cdot 10^{-3}$
$[C^{T,RR}_{ed}]_{2233}$ $-2\cdot 10^{-8}$ $-1\cdot 10^{-4}$ $-6\cdot 10^{-3}$
$[C^{T,RR}_{ed}]_{3311}$ $-2\cdot 10^{-3}$ $-1\cdot 10^{-1}$
$[C^{T,RR}_{ed}]_{3322}$ $-2\cdot 10^{-3}$ $-1\cdot 10^{-1}$
$[C^{T,RR}_{ed}]_{3333}$ $-4\cdot 10^{-7}$ $-2\cdot 10^{-3}$ $-1\cdot 10^{-1}$ $9\cdot 10^{-8}$
$[C^{T,RR}_{eu}]_{2211}$ $2\cdot 10^{-4}$ $-6\cdot 10^{-3}$
$[C^{T,RR}_{eu}]_{2222}$ $1\cdot 10^{-6}$ $2\cdot 10^{-4}$ $-6\cdot 10^{-3}$
$[C^{T,RR}_{eu}]_{3311}$ $4\cdot 10^{-3}$ $-1\cdot 10^{-1}$
$[C^{T,RR}_{eu}]_{3322}$ $2\cdot 10^{-5}$ $4\cdot 10^{-3}$ $-1\cdot 10^{-1}$
$[C^{S,RR}_{ed}]_{1133}$ $-2\cdot 10^{-4}$ $-3\cdot 10^{-8}$
$[C^{S,RR}_{ed}]_{2211}$ $-4\cdot 10^{-8}$ $-6\cdot 10^{-6}$
$[C^{S,RR}_{ed}]_{2222}$ $-4\cdot 10^{-8}$ $-6\cdot 10^{-6}$
$[C^{S,RR}_{ed}]_{2233}$ $-4\cdot 10^{-8}$ $-6\cdot 10^{-6}$
$[C^{S,RR}_{ed}]_{3311}$ $-7\cdot 10^{-7}$ $-10\cdot 10^{-5}$
$[C^{S,RR}_{ed}]_{3322}$ $-7\cdot 10^{-7}$ $-10\cdot 10^{-5}$
$[C^{S,RR}_{ed}]_{3333}$ $-7\cdot 10^{-7}$ $-10\cdot 10^{-5}$
$[C^{S,RR}_{eu}]_{2211}$ $-2\cdot 10^{-7}$ $1\cdot 10^{-5}$
$[C^{S,RR}_{eu}]_{2222}$ $-2\cdot 10^{-7}$ $1\cdot 10^{-5}$
$[C^{S,RR}_{eu}]_{3311}$ $-3\cdot 10^{-6}$ $2\cdot 10^{-4}$
$[C^{S,RR}_{eu}]_{3322}$ $-1\cdot 10^{-8}$ $-3\cdot 10^{-6}$ $2\cdot 10^{-4}$
$[C^{V,LR}_{ee}]_{1221}$ $-2\cdot 10^{-5}$
$[C^{V,LR}_{ee}]_{1331}$ $-2\cdot 10^{-5}$
$[C^{S,RR}_{ee}]_{1221}$ $6\cdot 10^{-4}$
$[C^{S,RR}_{ee}]_{1331}$ $1\cdot 10^{-2}$
$[C^{S,RR}_{ee}]_{1122}$ $-1\cdot 10^{-5}$
$[C^{S,RR}_{ee}]_{1133}$ $-2\cdot 10^{-4}$
@@
Table 6: At the 2-loop level, the operator mixing of the semileptonic operators in $5+3$ WET (1st column) onto dipole operators (2nd row) in $3+1$ and $5+3$ flavor WET is shown. Here the entries represent the quantity $\eta_{i}^{j}(\mu_{\rm low},\mu_{\rm ew})$ Entries below $10^{-8}$ are dropped.
$C_{j}(\mu_{low})$ stands for the WCs of the dipole and Weinberg operators that enter at tree-level to the EDM expressions. The $C_{i}(\mu_{ew})$ represent the all other WET operators which mix with $C_{j}(\mu_{low})$ via any of the mechanism discussed above. In the 4+1 and 5+3 theories we only show the operators that contain at least one heavy flavor fermion current. Running effects, in particular QCD running, are also important for light flavor operators, and they are fully accounted for in the master formulae that we provide in Section 5.
@@ 5 EDM Master Formulae in WET
$d_{X}={\rm abs}\left(\sum_{I\in\rm dipoles}\alpha_{I}^{X}(\mu_{ew}){C_{I}^{(5)}(\mu_{ew})\over[{\rm TeV}]}+\sum_{J\in\rm 4f}\alpha_{J}^{X}(\mu_{ew}){C_{J}^{(6)}(\mu_{ew})\over[{\rm{TeV^{2}}}]}\right)~{}e\,{\rm cm}\,,$ (115)
here $X=n,p$ or Hg, Xe, Ra and the indices $I$ and $J$ run over all contributing dipole and four-fermion operators in the $n_{f}=5+3$ $\Delta F=0$ WET WCs at the EW scale. The complete list of relevant operators can be found in Sec. 2. The coefficients $\alpha_{I}$ encode the long-distance effects due to matrix elements and RG running at the 2-loop level as described in the previous section. These coefficients have units of $e\,\rm cm\,{TeV}^{2}$ or $e\,\rm cm\,TeV$ for the dim-6 and dim-5 WCs, respectively. The WCs in (115) have $\rm{TeV}^{-2}$ and TeV units for dim-6 and five operators. The overall units of $d_{X}$ are $e$ cm. To obtain the EDMs in the $\rm{GeV}^{-1}$ units we need to divide the $\alpha_{I}$ coefficients by a conversion factor according to ${\rm GeV}^{-1}=6.52\times 10^{-14}e\,\textrm{cm}$.
We can write very similar expressions for the frequencies for the electron EDM. In this case, the $\alpha_{I}^{X}$ coefficients will have (mrad/s)$\rm{TeV}^{2}$ or (mrad/s)TeV units. The values for $\alpha_{I}$ coefficients at the EW scale and their theoretical uncertainties for the HfF precession frequency, and the neutron and Hg EDMs are given in Tab. 7, 10, and 8, 9, respectively. The master formulae for $d_{p}$, $\omega_{\rm ThO}$ and $\omega_{\rm YbF}$ can be found in appendix A. We do not provide explicit master formulae for $d_{\rm Xe}$ and $d_{\rm Ra}$, but they have been implemented in flavio.
@@ 5.1 Theoretical uncertainties
The theoretical uncertainties in the $\alpha_{I}^{X}$ coefficients can stem from various sources such as the SM parameters and the matrix elements. The latter arise from both nuclear theory and hadronization uncertainties and, for most species, is a dominant source. The uncertainties from atomic theory are in most cases better understood. As discussed in Section 3, the quoted hadronic and nuclear uncertainties provide at best a very rough estimate of the real theoretical uncertainty. In this respect, the field is less advanced compared to flavor physics, where, for several observables, lattice QCD matrix elements have been evaluated, or collider physics. Here we assume theoretical uncertainties to be Gaussian distributed. The estimate of the theoretical uncertainties is based on the random sampling of the input parameters and the evaluation of the standard deviation of the corresponding EDM observables.
5.2 Master formula for $\omega_{\text{HfF}}$
In this subsection, we provide a master formula for $\omega_{\text{HfF}}$ (mrad/s) for the electron EDM, see Tab. 7. Similar formulae for the $\omega_{\text{ThO}}$ and $\omega_{\text{YbF}}$ are given in App. A.

### v2

@@ 4.5 Numerical solution for the RGEs
$[C^{S,RR}_{eu}]_{1111}$ $5.8\cdot 10^{-8}$ $-2.9\cdot 10^{-7}$
$[C^{S,RR}_{ed}]_{1111}$ $-2.9\cdot 10^{-8}$ $-1.5\cdot 10^{-7}$
$[C^{S,RR}_{ed}]_{1122}$ $-2.9\cdot 10^{-8}$ $-3.1\cdot 10^{-6}$
@@
$[C^{T,RR}_{eu}]_{1111}$ $1.5\cdot 10^{-6}$ $-3.0\cdot 10^{-5}$ $1.4\cdot 10^{-4}$
$[C^{T,RR}_{ed}]_{1111}$ $-5.8\cdot 10^{-7}$ $-3.0\cdot 10^{-5}$ $-1.5\cdot 10^{-4}$
$[C^{T,RR}_{ed}]_{1122}$ $-5.8\cdot 10^{-7}$ $-3.0\cdot 10^{-5}$ $-3.0\cdot 10^{-3}$
$[C^{T,RR}_{eu}]_{1122}$ $7.4\cdot 10^{-2}$ $6.7\cdot 10^{-7}$ $-2.6\cdot 10^{-5}$
$[C^{S,RR}_{eu}]_{1122}$ $-1.5\cdot 10^{-4}$ $4.0\cdot 10^{-8}$
$[C^{T,RR}_{ed}]_{1133}$ $-1.1\cdot 10^{-1}$ $-3.4\cdot 10^{-7}$ $-2.6\cdot 10^{-5}$
$[C^{T,RR}_{ed}]_{2211}$ $-1.2\cdot 10^{-4}$ $-6.3\cdot 10^{-3}$
$[C^{T,RR}_{ed}]_{2222}$ $-1.2\cdot 10^{-4}$ $-6.3\cdot 10^{-3}$
$[C^{T,RR}_{ed}]_{2233}$ $-3.8\cdot 10^{-8}$ $-7.0\cdot 10^{-5}$ $-5.4\cdot 10^{-3}$
$[C^{T,RR}_{ed}]_{3311}$ $-2.0\cdot 10^{-3}$ $-1.1\cdot 10^{-1}$
$[C^{T,RR}_{ed}]_{3322}$ $-2.0\cdot 10^{-3}$ $-1.1\cdot 10^{-1}$
$[C^{T,RR}_{ed}]_{3333}$ $-6.4\cdot 10^{-7}$ $-1.2\cdot 10^{-3}$ $-9.0\cdot 10^{-2}$
$[C^{T,RR}_{eu}]_{2211}$ $3.1\cdot 10^{-4}$ $-6.3\cdot 10^{-3}$
$[C^{T,RR}_{eu}]_{2222}$ $1.1\cdot 10^{-6}$ $1.4\cdot 10^{-4}$ $-5.3\cdot 10^{-3}$
$[C^{T,RR}_{eu}]_{3311}$ $5.3\cdot 10^{-3}$ $-1.1\cdot 10^{-1}$
$[C^{T,RR}_{eu}]_{3322}$ $1.9\cdot 10^{-5}$ $2.3\cdot 10^{-3}$ $-9.0\cdot 10^{-2}$
$[C^{S,RR}_{ed}]_{1133}$ $-8.1\cdot 10^{-5}$ $-2.0\cdot 10^{-8}$
$[C^{S,RR}_{ed}]_{2211}$ $-3.6\cdot 10^{-8}$ $-6.0\cdot 10^{-6}$
$[C^{S,RR}_{ed}]_{2222}$ $-3.6\cdot 10^{-8}$ $-6.0\cdot 10^{-6}$
$[C^{S,RR}_{ed}]_{2233}$ $-4.1\cdot 10^{-6}$
$[C^{S,RR}_{ed}]_{3311}$ $-6.0\cdot 10^{-7}$ $-1.0\cdot 10^{-4}$
$[C^{S,RR}_{ed}]_{3322}$ $-6.0\cdot 10^{-7}$ $-1.0\cdot 10^{-4}$
$[C^{S,RR}_{ed}]_{3333}$ $-6.9\cdot 10^{-5}$
$[C^{S,RR}_{eu}]_{2211}$ $-2.6\cdot 10^{-7}$ $1.2\cdot 10^{-5}$
$[C^{S,RR}_{eu}]_{2222}$ $8.2\cdot 10^{-6}$
$[C^{S,RR}_{eu}]_{3311}$ $-4.4\cdot 10^{-6}$ $2.0\cdot 10^{-4}$
$[C^{S,RR}_{eu}]_{3322}$ $-1.6\cdot 10^{-8}$ $1.4\cdot 10^{-4}$
$[C^{V,LR}_{ee}]_{1221}$ $-2.0\cdot 10^{-4}$
$[C^{V,LR}_{ee}]_{1331}$ $-3.2\cdot 10^{-3}$
$[C^{S,RR}_{ee}]_{1221}$ $6.3\cdot 10^{-4}$
$[C^{S,RR}_{ee}]_{1331}$ $1.1\cdot 10^{-2}$
$[C^{S,RR}_{ee}]_{1122}$ $-1.3\cdot 10^{-5}$
$[C^{S,RR}_{ee}]_{1133}$ $-2.2\cdot 10^{-4}$
@@
Table 6: Operator mixing of the semileptonic operators in $5+3$ WET (1st column) onto dipole operators (2nd row) in $3+1$ and $5+3$ flavor WET is shown. Here the entries represent the quantity $\eta_{i}^{j}(\mu_{\rm low},\mu_{\rm ew})$, $\mu_{\rm ew}=91.1876$ GeV, $\mu_{\rm low}=2.0$ GeV . Entries below $10^{-8}$ are dropped. Note that for the charm and bottom dipoles, we stop running at $\mu=m_{b}=4.18$ GeV.
In Tabs. 4, 5 and 6, we show the numerical values of the quantity
for $\mu_{low}=2$ GeV and $\mu_{ew}=91$ GeV. In these tables, we focus mainly on heavy flavor operators, that do not generate EDMs at tree level. In this case, $C_{j}(\mu_{low})$ stands for the WCs of the dipole and Weinberg operators that enter at tree-level to the EDM expressions. The $C_{i}(\mu_{ew})$ represent the all other WET operators which mix with $C_{j}(\mu_{low})$ via any of the mechanism discussed above. In the 4+1 and 5+3 theories we only show the operators that contain at least one heavy flavor fermion current. Running effects, in particular QCD running, are also important for light flavor operators, and they are fully accounted for in the master formulae that we provide in Section 5.
@@ 5 EDM Master Formulae in WET
$d_{X}=\sum_{I\in\rm dipoles}\alpha_{I}^{X}(\mu_{ew}){C_{I}^{(5)}(\mu_{ew})\left[{\rm TeV^{-1}}\right]}+\sum_{J\in\rm 4f}\alpha_{J}^{X}(\mu_{ew}){C_{J}^{(6)}(\mu_{ew})\left[{\rm{TeV^{-2}}}\right]}\,,$ (115)
here $X=n,p$ or Hg, Xe, Ra and the indices $I$ and $J$ run over all contributing dipole and four-fermion operators in the $n_{f}=5+3$ $\Delta F=0$ WET at the EW scale. The complete list of relevant operators can be found in Sec. 2. The coefficients $\alpha_{I}$ encode the long-distance effects due to matrix elements and RG running at the 2-loop level as described in the previous section. These coefficients have units of $e\,\rm cm\,{TeV}^{2}$ or $e\,\rm cm\,TeV$ for the dim-6 and dim-5 WCs, respectively. The WCs in (115) have $\rm{TeV}^{-2}$ and TeV-1 units for dim-6 and dim-5 operators. This is indicated by the units in square brackets in Eq. (115). In this way, the overall units of $d_{X}$ are $e$ cm. To obtain the EDMs in the $\rm{GeV}^{-1}$ units we need to divide the $\alpha_{I}$ coefficients by a conversion factor according to ${\rm GeV}^{-1}=6.52\times 10^{-14}e\,\textrm{cm}$.
We can write very similar expressions for the frequencies for the electron EDM. In this case, the $\alpha_{I}^{X}$ coefficients will have (mrad/s)$\rm{TeV}^{2}$ or (mrad/s)TeV units. The values for $\alpha_{I}$ coefficients at the EW scale and their theoretical uncertainties for the HfF precession frequency, the Hg and neutron EDMs are given in Tab. 7, 8, 9 and 10. The master formulae for $d_{p}$, $\omega_{\rm ThO}$ and $\omega_{\rm YbF}$ can be found in appendix A. We do not provide explicit master formulae for $d_{\rm Xe}$ and $d_{\rm Ra}$, but they have been implemented in flavio.
@@ 5.1 Theoretical uncertainties
The theoretical uncertainties in the $\alpha_{I}^{X}$ coefficients can stem from various sources such as the SM parameters and the matrix elements. The latter arise from both nuclear theory and hadronization uncertainties and, for most species, they are the dominant source. The uncertainties from atomic theory are in most cases better understood. As discussed in Section 3, the quoted hadronic and nuclear uncertainties provide at best a very rough estimate of the real theoretical uncertainty. In this respect, the field is less advanced compared to flavor physics, where, for several observables, lattice QCD matrix elements have been evaluated, or collider physics. Here we assume theoretical uncertainties to be Gaussian distributed. The estimate of the theoretical uncertainties is based on the random sampling of the input parameters and the evaluation of the standard deviation of the corresponding EDM observables.
5.2 $\omega_{\text{HfF}}$
In this subsection, we provide a master formula for $\omega_{\text{HfF}}$ (mrad/s), which is sensitive to the electron EDM, see Tab. 7. Similar formulae for the $\omega_{\text{ThO}}$ and $\omega_{\text{YbF}}$ are given in App. A.